OpenStudy (anonymous):

a sled travels 15 meters down a slope inclined at 30° with the horizontal. what is the horizontal displacement of the sled?

OpenStudy (ivancsc1996):

So if you look a it, it is a right triangle. You just need to solve the sin(30) for S. \[\sin(30)=S/15m \rightarrow S=15msin(30)=7.5m\]

OpenStudy (anonymous):

The horizontal displacement \[\cos30=S/15\] \[S=15 \cos30 = 15\times \sqrt{3}/2=12.99 m\]